### v2

Abstract
Difference-in-differences (DID) is commonly used to estimate treatment effects but is infeasible in settings where data are unpoolable due to privacy concerns or legal restrictions on data sharing, particularly across jurisdictions. In this study, we identify and relax the assumption of data poolability in DID estimation. We propose an innovative approach to estimate DID with unpoolable data (UN-DID) which can accommodate covariates, multiple groups, and staggered adoption. Through analytical proofs and Monte Carlo simulations, we show that UN-DID and conventional DID estimates of the average treatment effect and standard errors are equal and unbiased in settings without covariates. With covariates, both methods produce estimates that are unbiased, equivalent, and converge to the true value. The estimates differ slightly but the statistical inference and substantive conclusions remain the same. Two empirical examples with real-world data further underscore UN-DID’s utility. The UN-DID method allows the estimation of cross-jurisdictional treatment effects with unpoolable data, enabling better counterfactuals to be used and new research questions to be answered.
Keywords: difference-in-differences, treatment effects, siloed data, unpooled data, impact evaluation
JEL Codes: C10, C12, C21, D04, I18, I38, Z18
1 Introduction
Difference-in-differences (DID) is a widely used method to evaluate the impacts of policies and interventions. It is used to estimate both intended treatment effects and unintended consequences. The source of identifying variation in DID analyses is often a policy change that occurs in some political jurisdictions (e.g., states, countries) but not in others. Treated observations come from jurisdictions or units in which the intervention occurs, while control or untreated observations come from jurisdictions that do not experience that change (at least not at the same time). DID can be used to estimate both the intent-to-treat (ITT) effect and the average treatment effect on the treated (ATT). In the absence of randomized treatment assignment, a widely documented set of assumptions support the unbiased estimation of these effects. These assumptions include parallel pre-intervention trends, no anticipation, no staggered adoption, homogeneous treatment effects, and a single isolated treatment (Roth et al., 2022; Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a; Callaway and Sant’Anna, 2021; De Chaisemartin and D’Haultfœuille, 2020b). The strong correlation between treatment assignment and jurisdiction implies an additional assumption underlying the use of DID estimation that has not yet been made explicit: that data from the treatment group and the control group can be combined together for analyses.
This paper develops the DID for unpoolable data (UN-DID) method, enabling the estimation of treatment effects using DID when the data pooling assumption is violated. This occurs when data come from the same jurisdictional level as the policy variation, and may not be combined on the same computer or server, usually due to privacy concerns or legal restrictions on data sharing. The data are therefore siloed at the level of treatment, and traditional DID methods cannot be used to estimate the ATT. Examples include health care data from different insurers (e.g., Canadian provinces (Health Data Research Network Canada, 2021), private and public insurers, all-payer claims databases across states (Kranz et al., 2023)), restricted-use micro-data held in secure environments (e.g., vital statistics, income tax, business data held in Statistics Canada’s or the US Census Bureau’s Research Data Centres (Canadian Research Data Centre Network, 2024; United States Census Bureau, 2024)), and electronic health records federated across countries (e.g., the European Health Data Space (European Commission, 2024)). Sometimes data are considered unpoolable because the datasets are not sufficiently harmonized. We do not address that issue here - UN-DID requires that data that are physically unpoolable be sufficiently harmonized (either de facto or after cross-jurisdictional collaboration). That said, UN-DID is able to accommodate datasets from different jurisdictions that measure the same underlying concepts, but code or categorize variables slightly differently (e.g., income in different bins). UN-DID enables the estimation of treatment effects when policies and interventions vary at the level of data silos (county, state, country, etc.), opening opportunities to estimate less-biased treatment effects using better controls and to answer new research questions.
When the ability to pool data across treatment and control jurisdictions is limited, the choice of appropriate counterfactuals is necessarily constrained. Researchers may resort to estimating first-differences (Fischer et al., 2021) or using interrupted time series (Tamblyn, 2001; Bernal et al., 2017; Farin et al., 2024), which only provide unbiased treatment effect estimates if no other changes coincide with the intervention of interest. More convincing treatment effect estimates are obtained leveraging within-jurisdiction controls (Rudoler et al., 2023; Strumpf et al., 2017; Li et al., 2014; Girma and Görg, 2007) or individual fixed effects (Dumont et al., 2008), both of which may still suffer from selection on unobservables. The impacts of national policies that take effect everywhere at the same time can sometimes be identified using pre-policy variation (Finkelstein, 2007; Duflo, 2001), but in other cases, including nationwide cannabis legalization in Canada in 2018, rigorous policy evaluations using external controls remain extremely limited or non-existent.
If policy designs permit the use of RD or IV to estimate local average treatment effects (Almond et al., 2010; Hutcheon et al., 2020), the identifying variation may be convincing, but the generalizability of the estimate narrow by definition. Meta-analysis can be used with unpoolable data, and is an effective way to leverage sample size by combining silo-specific treatment effect estimates into a single parameter (Xiong et al., 2023). For example, the Canadian Network for Observational Drug Effect Studies (CNODES) routinely employs meta-analytical techniques to more comprehensively examine drug safety using siloed provincial or national data (Suissa et al., 2012; Filion et al., 2014; Durand et al., 2021). However, meta-analysis cannot be used to estimate treatment effects of an intervention in one jurisdiction using another unexposed jurisdiction as a control. UN-DID will allow researchers to estimate treatment effects using better counterfactuals, leveraging data from non-treated jurisdictions and limiting selection bias while maintaining generalizability.
Data poolability limitations can also constrain which research questions get asked, specifically which outcomes are examined in evaluations of policy interventions. National and international surveys include multiple states/provinces or countries, so DID models are easily estimated with treated and control units in the same dataset (Baker et al., 2008; Charters et al., 2013; Dube, 2019). The range of outcomes available in a survey, however, are necessarily limited with respect to content, detail, and method of reporting (self-report vs. administrative data). In the absence of a method like UN-DID to allow treatment effect estimation across separate datasets, the questions that have been asked and answered have been a function of the (pooled) datasets that are available. This constraint has become even more problematic in recent years, precisely because large, detailed, and high-quality data are increasingly available for research. All-payer health insurance claims, electronic medical records, detailed individual- and firm-level tax filings, genetic data, and immigration records offer comprehensiveness, accuracy and detail that allow researchers to ask and answer new questions. However, these data are generally accessible only in secure research data centers and are not shareable or poolable across jurisdictions. By enabling treatment effect estimation using siloed data, UN-DID will open up opportunities to answer novel research questions and exploit new sources of identifying variation using separate datasets across jurisdictions.
The UN-DID method allows the estimation of treatment effects with unpoolable data. This has important implications for research across the fields of applied economics, and across a range of disciplines including public policy, epidemiology, health research, environmental science, political science, sociology, and others. With UN-DID, new research questions can be answered and better counterfactuals can be used to reduce bias when estimating treatment effects. While researchers with treatment and control observations in separate datasets could already calculate the double-difference “by hand”, our formal results justify this approach and extend beyond the 2$\times$2 case. Furthermore, UN-DID enables the estimation of appropriate standard errors and therefore proper inference. We are developing UN-DID software packages and a user guide that will follow as a separate publication. These will enable UN-DID to be implemented across secure data silos, even when local data analysts are not skilled econometricians.
The intuition behind UN-DID is based on non-parametric DID estimation where conditional mean outcomes are estimated for each of four treatment*timing groups: pre-intervention treatment, pre-intervention control, post-intervention treatment and post-intervention control. The post-pre first difference can then be calculated for each treatment and control group, and then the DID estimate is simply the difference between those first differences. UN-DID estimates first-differences within jurisdictional silos and provides a method to combine these estimates into a treatment effect estimate and to test hypotheses. We provide formal results which show that this method accurately estimates the ATT. UN-DID also estimates the parameters necessary to evaluate parallel pre-intervention trends.
In this paper, we begin by reviewing the conventional DID identifying assumptions and formally introduce the poolability assumption, that data can be combined across treatment and control groups. We start with a one treatment, one control, and several time periods (2$\times$×T) setting. We propose a method for estimating the ATT when the poolability assumption is violated, which we call UN-DID (DID for unpoolable data). Because researchers commonly rely on regression-based tools for estimating treatment effects, we introduce the UN-DID method as a regression-based tool for this purpose. Through analytical proofs, we first demonstrate the equivalence of estimates between conventional DID and UN-DID without covariates. We then relax the strong parallel trends assumption and add the assumption of common causal covariates (each covariate has a common slope across silos). In this setting, we show the equivalence of conventional DID and UN-DID with time-invariant covariates, and in larger samples with time-varying covariates. Lastly, we turn to the many group many time period (G×$\times$×T) setting and relax the no staggered adoption assumption. We show how the UN-DID method works in the staggered adoption setting in the spirit of Callaway and Sant’Anna (2021). Throughout, we also demonstrate the equivalence of conventional standard errors between conventional DID and UN-DID. Cluster robust inference is possible as described in work in progress by the authors of this paper.
With analytical results in hand, we conduct Monte Carlo experiments with simulated panel data designed to mimic populations in Canada’s two largest provinces over 10 years (2$\times$10). We consider data generating processes (DGPs) with and without treatment effects and with no covariates, with a time-fixed and with a time-varying covariate. We consider equal and unequal sample sizes between the two provinces, and vary the sample size from 500 to 50,000 observations. To assess unbiasedness, we plot kernel densities of UN-DID and conventional DID ATT estimates against the known ATT value from the DGP. To assess the asymptotic performance of UN-DID and the conventional DID, we compare estimates with the true values of the ATT and standard errors against to the known true values as the sample size increases. Specifically, we calculate the mean squared error between the ATTs and the SEs for each pair: UN-DID vs. truth, convential DID vs. truth, and UN-DID vs. conventional DID. Lastly, to assess UN-DID’s performance and to highlight its utility, we include two empirical examples: one with common treatment timing and one with staggered adoption. Here, real poolable data are treated as if they were unpoolable across treated and control groups, allowing us to compare the estimated ATTs and SEs from UN-DID and conventional DID with real-world data.
In the simulation study, we demonstrate that UN-DID and conventional DID are both unbiased and asymptotically consistent without covariates. With a time-invariant covariate, the two ATTs are unbiased and converge to the true value; the SEs are equivalent but not numerically equal. With a time-varying covariate, the UN-DID and conventional DID ATTs and SEs are unbiased, equivalent, and converge to the true value. The above results hold with both equal and unequal sample sizes across silos. These results hold in our two empirical examples using real-world data. Without covariates, UN-DID’s ATT and SE estimates are exactly the same as conventional DID. With covariates, the estimates differ slightly but the statistical inference and substantive conclusions remain the same. As the number of covariates and/or data silos increases, the more flexible UN-DID model may differ more from conventional DID which constrains covariates to have the same slopes across silos. These results reinforce the theoretical and simulation study findings, that the UN-DID procedure estimates either identical or very similar treatment effects compared to the conventional method.
This paper contributes to the dynamic literature on DID methods. In the past five years, an important series of papers have looked under the hood of the conventional DID estimator, exploring what is actually being estimated and what can go wrong when each of the identifying assumptions is violated (Roth et al., 2022; Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a; Callaway and Sant’Anna, 2021; De Chaisemartin and D’Haultfœuille, 2020b; Goodman-Bacon, 2021). They have developed methods to identify the ATT in each case, changing the way that DID analyses are conducted. We highlight the heretofore implicit assumption of data poolabilty and provide a solution to estimate the ATT when this assumption does not hold. We assess UN-DID’s performance and demonstrate its equivalence to conventional DID.
We evidently aim to contribute to the large literature that uses DID to estimate treatment effects of policy changes and other interventions, both in economics and across a wide range of disciplines. In this paper, we apply UN-DID to two published policy evaluations that used conventional DID methods (Sabia et al., 2017; Conley and Taber, 2011). Our primary goal is to compare the estimates from the two methods, with a secondary goal of highlighting how UN-DID could be used in empirical analyses. The forthcoming software package user guide will use UN-DID to evaluate a policy change using real-world siloed data, where a conventional DID analysis is impossible and no ATT estimates exist for comparison.
The paper proceeds as follows: Section 2 presents the theoretical framework and the identifying assumptions underlying conventional DID. Section 3 demonstrates the equivalence of the ATT estimated with UN-DID and with conventional DID when the strong parallel trends assumption holds, describes the basic Monte Carlo design, and presents simulation results for the cases without covariates. Section 4 demonstrates the equivalence UN-DID and conventional DID under the weaker assumption of conditional parallel trends and of common causal covariates, expands the Monte Carlo design to include time-fixed and time-varying covariates, and presents results. Section 5 presents UN-DID with staggered adoption. Section 6 describes a simplified method of aggregating estimates from multiple silos. Section 7 compares UN-DID and conventional DID ATT estimates for two empirical examples using real-world data and Section 8 concludes.
2 Theoretical Framework
Let us assume that we have data for two groups, the treatment group which received the intervention or treatment and the control group which did not. We also assume data for several calendar years, $t=1,2,.....,R$ and that the treatment was implemented at year $k$. All calendar years before $k$ ($t<k$) are pooled together as the pre-intervention period. Similarly, all calendar years from $k$ onward ($t\geq k$) are pooled together as the post-intervention period. It is common practice to estimate DID using micro-level data, where numerous observations from individual subjects are available within each of the relevant cells.
The basic idea of DID estimation is to compare the difference in outcomes before and after the treatment between groups that received the treatment and groups that did not (Bertrand et al., 2004). The change in outcomes for the control group is used as the unobserved counterfactual of the treated group. Card and Krueger (1993) first used the conventional DID to estimate the effect of an increase in minimum wage on unemployment in the American state of Pennsylvania. The conventional way of estimating the ATT uses the following regression:
$Y^{S}_{i,t}=\beta_{0}+\beta_{1}D^{S}_{i}+\beta_{2}P^{S}_{t}+\beta_{3}P^{S}_{t}*D^{S}_{i}+\epsilon^{S}_{i,t}.$ (1)
Here, $D^{S}_{i}$ is a dummy variable that takes on a value of 1 if individual $i$ is in the treatment group, and 0 otherwise. $P^{S}_{t}$ is a dummy variable that takes on a value of 1 if the observation is in the post-intervention period, and 0 otherwise. $S$ implies that we are combining the dataset from both the treated group $T$ and the control group $C$ into a common dataset ($S=T+C$). Under very strong assumptions, $\hat{\beta_{3}}$ (the coefficient of the interaction term between $D^{S}_{i}$ and $P^{S}_{t}$) identifies the ATT (Roth et al., 2022). To simplify notations, we will not use the $S$ superscripts to denote a combined dataset for the rest of the paper.
Assumption 1: Treatment is Binary
This implies individual $i$ can be either treated or not treated at time $t$. There are no variations in treatment intensity.
$D_{i}=\begin{cases}1&\mbox{if individual i is treated at time t}.\\ 0&\mbox{if individual i is not treated at time t}.\\ \end{cases}$
Assumption 2: Strong Parallel Tends
$\begin{split}&\biggr{[}E[Y_{i}(0)|D_{i}=1,P_{t}=1]-E[Y_{i}(0)|D_{i}=1,P_{t}=0]\biggr{]}\\ &=\ \biggr{[}E[Y_{i}(0)|D_{i}=0,P_{t}=1]-E[Y_{i}(0)|D_{i}=0,P_{t}=0]\biggr{]}.\end{split}$ (2)
Here, the $Y_{i}(0)$ are the untreated potential outcomes for the relevant group and period. Strong parallel trends imply that the evolution of outcome between treated and control groups immediately before treatment are the same. This ensures that the selection bias is 0 (Roth et al., 2022). When strong parallel trends hold, the ATT is shown in Equation (3). Refer to Callaway and Sant’Anna (2021) for a simple proof.
$\footnotesize\begin{gathered}ATT=\biggl{[}E[Y_{i}|D_{i}=1,P_{t}=1]-E[Y_{i}|D_{i}=1,P_{t}=0]\biggr{]}-\biggl{[}E[Y_{i}|D_{i}=0,P_{t}=1]-E[Y_{i}|D_{i}=0,P_{t}=0]\biggr{]}.\end{gathered}$ (3)
Here, $E[Y_{i}|D_{i}=1,P_{t}=1]$ is the expected outcome of the treated group in the post-intervention period and $E[Y_{i}|D_{i}=1,P_{t}=0]$ is the expected outcome of the treated group in the pre-intervention period. Similarly, $E[Y_{i}|D_{i}=0,P_{t}=1]$ is the expected outcome of the control group in the post-intervention period and $E[Y_{i}|D_{i}=0,P_{t}=0]$ is the expected outcome of the control group in the pre-intervention period.
Assumption 3: No Anticipation
No anticipation implies that treated units do not change behavior before treatment occurs (Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a). So, the treated potential outcome is equal to the untreated potential outcome for all units in the treated group in the pre-intervention period. Violation of no anticipation can also lead to deviations in parallel trends before treatment. Here, $Y_{i}(t)$ is the treated potential outcome of individual $i$ at calendar year $t$.
$\begin{gathered}\biggl{[}E[Y_{i}(t)|D_{i}=1,P_{t}=0]=E[Y_{i}(0)|D_{i}=1,P_{t}=0]\biggr{]}\mbox{\quad{a.s.} for all}\;\;t<k.\end{gathered}$ (4)
Assumption 4: Homogeneous Treatment Effect Homogeneous treatment effect implies that all treated units have the same treatment effect across both time and individuals. Formally, it means that the difference in potential outcomes for treated units is the same for all time periods after treatment.
$\begin{gathered}\biggl{[}E[Y_{i}(t)|D_{i}=1,P_{t}=1]-E[Y_{i}(0)|D_{i}=1,P_{t}=1]\biggr{]}\\ =\biggl{[}E[Y_{j}(t)|D_{j}=1,P_{t}=1]-E[Y_{j}(0)|D_{j}=1,P_{t}=1]\biggr{]}\mbox{\quad{a.s.} for all}\;\;i\neq j.\end{gathered}$ (5)
Assumption 5: No Staggered Adoption
No staggered adoption implies treatment occurs only once (Callaway and Sant’Anna, 2021; De Chaisemartin and d’Haultfoeuille, 2020a). Formally, it means that all units in the treated group are treated at calendar year $k$. Throughout this paper, we maintain the above assumptions unless stated otherwise.
The papers cited above have contributed to a rich recent literature investigating what occurs when each of the aforementioned assumptions are violated. In this paper, we introduce a new assumption that has been implied in previous literature but has not been explicitly explored or discussed in the context of DID.
Assumption 6: Data are Poolable
Data from the treatment and control groups are available on a single server and can be combined together for analyses. $Y_{T}$ and $Y_{C}$ are two matrices containing the outcome variable for group $T$ and $C$, respectively, and $N_{T}$ and $N_{C}$ the total number of observations in the matrices, respectively.
$Y_{T}=\begin{bmatrix}y_{T1}\\ y_{T2}\\ \vdots\\ y_{TN_{T}}\end{bmatrix},Y_{C}=\begin{bmatrix}y_{C1}\\ y_{C2}\\ \vdots\\ y_{CN_{C}}\end{bmatrix}$
Under the poolability assumption, the two matrices can be stacked together into a common matrix Y, which is not feasible when there are legal restrictions preventing them from being stacked together (data are siloed or unpoolable):
$Y=\begin{bmatrix}Y_{T}\\ Y_{C}\end{bmatrix}=\begin{bmatrix}y_{T1}\\ y_{T2}\\ \vdots\\ y_{TN_{T}}\\ y_{C1}\\ y_{C2}\\ \vdots\\ y_{CN_{C}}\end{bmatrix}$
For settings where Assumption 6 is violated and the data are not poolable, we introduce the UN-DID method to estimate the ATT. In such cases, when we access the data from the treated silo, we have no knowledge of the data from the control silo, $Y_{C}$. Similarly, once we have access to the data from the untreated silo, we have no information about the data from the treated silo, or $Y_{T}$. The conventional DID method becomes impractical if siloed datasets cannot be combined for regression analysis.
In this case, we can visit each silo (physically or virtually) and run regressions shown in equations (6) and (7) for the treated and untreated silos, respectively. In the simplest case, we assume that there are only two silos: the treated silo ($T$) and the untreated silo ($C$). We also assume that there are no covariates. Here, $Y^{T}_{i,t}$ is the outcome for an individual from the silo that is treated at time $t$ and $Y^{C}_{i,t}$ is the outcome of an individual from the silo that is untreated at time $t$. $post_{t}^{T}$ is a dummy variable that takes on a value of 1 when the treated observation is in the post-intervention period, and 0 otherwise. Similarly, $post_{t}^{C}$ is a dummy variable that takes on a value of 1 when the untreated observation is in the post-intervention period, and 0 otherwise. $pre_{t}^{T}$ is a dummy variable that takes on a value of 1 if the treated observation is in the pre-intervention period, hence $pre_{t}^{T}=1-post_{t}^{T}$. Similarly, $pre_{t}^{C}$ is a dummy variable that takes on a value of 1 if the untreated observation is in the pre-intervention period, hence $pre_{t}^{C}=1-post_{t}^{C}$. Note that the regressions in equations (6) and (7) do not include a constant. Therefore, none of the variables are dropped because of multicollinearity.
[…]
These regressions are silo-specific. Because data are siloed, the treated regression does not contain any data from the untreated silo and the untreated regression does not contain any data from the treated silo. Here, $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ is the estimate of the ATT, as shown in Equation (8). The associated standard error of the ATT is estimated as the square root of the sum of the variances used to estimate the ATT estimated from the above UN-DID regressions minus twice the covariances between the relevant coefficients. This is shown in Equation (9).
$\widehat{ATT}=(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda_{1}}^{C}).$ (8)
$\small\widehat{SE}(\widehat{ATT})=\sqrt{\widehat{SE}(\widehat{\lambda}_{1}^{T})^{2}+\widehat{SE}(\widehat{\lambda_{2}}^{T})^{2}+\widehat{SE}(\widehat{\lambda}_{1}^{C})^{2}+\widehat{SE}(\widehat{\lambda}_{2}^{C})^{2}-2\widehat{Cov}(\widehat{\lambda}_{1}^{T},\widehat{\lambda}_{2}^{T})-2\widehat{Cov}(\widehat{\lambda}_{1}^{C},\widehat{\lambda}_{2}^{C})}.$ (9)
2.1 Cluster-Robust Inference
In this paper, we assume that the error terms are independent. Accordingly, we can estimate standard errors that are robust to heteroskedasticity but not robust to clustering. It is straightforward to extend the procedures that we discuss to clustering at the sub-silo level. For instance, one could cluster at the person-level when using panel data, or at the city level if the silos are state level. However, researchers will typically cluster at the level of the policy change (Bertrand et al., 2004; MacKinnon et al., 2023), which we have assumed is by silo. Silo level clustering is challenging when models like those in equations (6) and (7) are also estimated at the silo level. Conventional methods of cluster-robust inference will not work with silo-specific data because there is only one cluster in each dataset. With only one cluster the variance matrix is rank deficient and cannot be computed. Work in progress by the authors of this paper aims to extended these methods to allow for silo-level clustering using alternative inference procedures.
3 UN-DID When Strong Parallel Trends Hold
In this section, we assume that the strong parallel trends assumption is plausible between the treated and untreated silos. We also assume that no anticipation, homogeneous treatment effects, and no staggered adoption hold. Under these assumptions, we show that the UN-DID and the conventional DID recover the same estimate of the ATT, both analytically and through a Monte Carlo simulation study.
3.1 Equivalence of the Estimated ATT Between Conventional and UN-DID Methods with No Covariates
Conventional Regression
Under assumptions 1 to 6, $\hat{\beta_{3}}$ in the regression shown in Equation (10) is the estimate of the ATT (Roth et al., 2022). Note that the conventional regression can be used only if the data poolability assumption holds. For simplicity of notation, we assign all calendar years in the pre-intervention period as $t=0$ and assign all calendar years in the post-intervention period as $t=1$.
$Y_{i,t}=\beta_{0}+\beta_{1}D_{i}+\beta_{2}P_{t}+\beta_{3}P_{t}*D_{i}+\epsilon_{i,t}.$ (10)
Now we will show that $\hat{\beta_{3}}$ is the sample analogue of Equation (3), or the estimate of the ATT. According to the Frisch–Waugh–Lowell (FWL) theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (12). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from the regression shown in Equation (11).
$P_{t}*D_{i}=\alpha_{0}+\alpha_{1}P_{t}+\alpha_{2}D_{i}+u_{i,t}.$ (11)
$Y_{i,t}=\widehat{\beta_{3}}(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})+u^{\prime}_{i,t}.$ (12)
From Equation (12), we compute the residuals ($P_{t}*D_{i}-\widehat{P_{t}*D_{i}}$) for each observation. For a treated observation in the pre-intervention period, the residual is given by $-\hat{\alpha_{0}}-\hat{\alpha_{2}}$. Conversely, for a treated observation in the post-intervention period, the residual becomes $1-\hat{\alpha_{0}}-\hat{\alpha_{1}}-\hat{\alpha_{2}}$. Similarly, an untreated observation in the pre-intervention period has a residual of $-\hat{\alpha_{0}}$, while an untreated observation in the post-intervention period yields a residual of $-\hat{\alpha_{0}}-\hat{\alpha_{1}}$. The residuals for all individuals in their respective cohort are the same, as shown by De Chaisemartin and d’Haultfoeuille (2020a). We then proceed to derive $\widehat{\beta}_{3}$ from Equation (12) by using the following OLS formula:
$\begin{split}\widehat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)}{\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)^{2}}.\end{split}$ (13)
We can substitute the computed residuals into the numerator and simplify the expression from Equation (13) as outlined below:
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)+\\ &\sum_{i\in t=0,C}Y_{i,1}^{C}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}\right).\end{split}$
Likewise, by substituting the residuals and simplifying, the denominator of Equation (13) can be expressed as
$\begin{split}A=\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)^{2}+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}\right)^{2}.\end{split}$
The denominator is the sum of squared residuals from the regression in Equation (11). Combining the numerator and the denominator, $\widehat{\beta}_{3}$ can be written as
$\begin{split}\widehat{\beta}_{3}=&\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)\sum_{i\in t=1,T}Y_{i,1}^{T}}{A}-\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)\sum_{i\in t=0,T}Y_{i,0}^{T}}{A}\\ &-\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}\right)\sum_{i\in t=1,C}Y_{i,1}^{C}}{A}+\frac{\left(-\widehat{\alpha}_{0}\right)\sum_{i\in t=0,C}Y_{i,0}^{C}}{A}.\end{split}$ (14)
In Appendix A, we show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group and the post-intervention period ($N_{1,1})$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group in the pre-intervention period ($N_{1,0}$). Similarly, $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the control group in the post-intervention period $(N_{0,1})$ and $\frac{\left(-\widehat{\alpha}_{0}\right)}{A}$ is reciprocal of the number of observations in the control group in the pre-intervention period ($N_{0,0})$. Therefore, Equation (14) can be further simplified as
$\begin{split}\hat{\beta}_{3}=\frac{1}{N_{1,1}}\sum_{i\in t=1,T}Y_{i,1}^{T}-\frac{1}{N_{1,0}}\sum_{i\in t=0,T}Y_{i,0}^{T}-\frac{1}{N_{0,1}}\sum_{i\in t=1,C}Y_{i,1}^{C}+\frac{1}{N_{0,0}}\sum_{i\in t=0,C}Y_{i,0}^{C}\end{split}$
$\begin{split}\Rightarrow\hat{\beta}_{3}=&\left(\overline{Y}^{T}_{i,1}-\overline{Y}^{T}_{i,0}\right)-\left(\overline{Y}^{C}_{i,1}-\overline{Y}^{C}_{i,0}\right).\end{split}$ (15)
In simpler notation, Equation (15) can be rewritten as
$\begin{split}\hat{\beta}_{3}=&\left(\overline{Y}_{D_{i}=1,P_{t}=1}-\overline{Y}_{D_{i}=1,P_{t}=0}\right)-\left(\overline{Y}_{D_{i}=0,P_{t}=1}-\overline{Y}_{D_{i}=0,P_{t}=0}\right).\end{split}$ (16)
Equation (16) is the sample analogue of the ATT under strong parallel trends shown in Equation (3).
UN-DID Regression
In this subsection, we hold that Assumption 6 is violated (data are no longer poolable), while assumptions 1 to 5 hold. Under the violation of Assumption 6, the conventional regression can no longer be used to estimate the ATT. Under these assumptions, we prove that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ derived from the UN-DID regressions can recover the estimate of the ATT. Provided that the data samples used for both approaches are the same, we demonstrate that the ATT obtained through the UN-DID regression is numerically equal to the conventional ATT estimate. Equations (6) and (7) can be rewritten in the following way:
$\;Y^{j}_{i,t}=\lambda_{1}^{j}pre_{t}^{j}+\lambda_{2}^{j}post_{t}^{j}+\nu^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;j=\{T,C\}.$ (17)
$\equiv\;Y^{j}_{i,t}=\gamma_{0}^{j}+\gamma_{1}^{j}post_{t}^{j}+\nu^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;j=\{T,C\}.$ (18)
It can be shown that Equation (17) is equivalent to Equation (18).
Proof: We know that $pre_{t}^{j}=(1-post_{t}^{j})$. Substituting this into Equation (17) gives
$Y^{j}_{i,t}=\lambda_{1}^{j}(1-post_{t}^{j})+\lambda_{2}^{j}post_{t}^{j}+\nu^{j}_{i,t}$
$\Rightarrow{Y^{j}_{i,t}}=\lambda_{1}^{j}+(\lambda_{2}^{j}-\lambda_{1}^{j})\text{post}_{t}^{j}+\nu^{j}_{i,t}.$ (19)
Comparing equations (18) and (19), we can see that $\lambda_{1}^{j}=\gamma_{0}^{j}$ and $(\lambda_{2}^{j}-\lambda_{1}^{j})=\lambda_{2}^{j}$. Now we will prove that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})=\hat{\gamma}_{1}^{T}-\hat{\gamma}_{1}^{C}$ is the sample analogue of the ATT shown in Equation (3). In Equation (18), the coefficient of interest is $\hat{\gamma}^{j}_{1}$. According to the FWL theorem, $\hat{\gamma^{j}_{1}}$ will be the same as the coefficient of $(post_{t}^{j}-\widehat{post}_{t}^{j})$ from the regression shown in Equation (21). Here, $(post_{t}^{j}-\widehat{post}t^{j})$ are the residuals from the regression shown in Equation (20).
$post_{t}^{j}=\eta_{0}^{j}+\omega^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;\;\;\eta_{0}^{j}=\overline{post}_{t}^{j}.$ (20)
$Y_{i,t}^{j}=\widehat{\gamma_{1}^{j}}(post^{j}_{t}-\widehat{post^{j}_{t}})+\nu^{j^{\prime}}_{i,t}.$ (21)
Similar to in the preceding subsection, we calculate the residuals $(post_{t}^{j}-\widehat{post_{t}^{j}})$ for each observation using Equation (20). For an observation in group j during the pre-intervention period, the residual is given by $(post^{j}_{t}-\widehat{post^{j}_{t}})=-\widehat{\eta_{0}^{j}}$. Likewise, for an observation in group j during the post-intervention period, the residual becomes $(post^{j}_{t}-\widehat{post^{j}_{t}})=1-\widehat{\eta_{0}^{j}}$. Following this, we proceed to obtain $\widehat{\gamma}_{1}^{j}$ from Equation (21) using the OLS formula:
$\begin{split}\widehat{\gamma}_{1}^{i}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\widehat{\eta}_{0}^{j}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\widehat{\eta}_{0}^{j}\right)}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{1-\widehat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\widehat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=0}Y_{i,0},\end{split}$
$\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.\quad$ (22)
Here, $B=\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}$. Following the proof in Appendix A, $\frac{1-\widehat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group $j$ in the post-intervention period and $\frac{\widehat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group j for the pre-intervention period. Therefore, Equation (22) can be rewritten as:
$\widehat{\gamma}_{1}^{j}=\frac{1}{N_{j,1}}\sum_{i\in t=1}Y_{i,1}-\frac{1}{N_{j,0}}\sum_{i\in t=0}Y_{i,0}$
$\Rightarrow\widehat{\gamma}_{1}^{j}=\left(\overline{Y}^{j}_{i,1}-\overline{Y}^{j}_{i,0}\right).$ (23)
From Equation (23), we can take $\widehat{\gamma}^{T}_{1}$ and $\widehat{\gamma}^{C}_{1}$ from the treated and untreated regressions, respectively, and then take the difference between them. Equation (24) shows that the difference is the sample analogue of the ATT shown in Equation (3). If we compare equations (16) and (25), we can see that the conventional DID estimates and the UN-DID estimates provide numerically equal estimates of the ATT. Note that the conventional regression is not possible with data that are siloed. In a hypothetical scenario, we can combine data from different silos together and run both the conventional and the UN-DID regressions on the data. In this case, both methods can identify the ATT and give equivalent estimates:
$\begin{split}\widehat{\gamma}_{1}^{T}-\widehat{\gamma}_{1}^{C}=\left(\overline{Y}^{T}_{i,1}-\overline{Y}^{T}_{i,0}\right)-\left(\overline{Y}^{C}_{i,1}-\overline{Y}^{C}_{i,0}\right)\end{split}$
$\begin{split}\Rightarrow(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})=\left(\overline{Y}^{T}_{i,1}-\overline{Y}^{T}_{i,0}\right)-\left(\overline{Y}^{C}_{i,1}-\overline{Y}^{C}_{i,0}\right).\end{split}$ (24)
In simpler notation, Equation (24) can be rewritten as
$\begin{split}(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})=&\left(\overline{Y}_{D_{i}=1,P_{t}=1}-\overline{Y}_{D_{i}=1,P_{t}=0}\right)-\left(\overline{Y}_{D_{i}=0,P_{t}=1}-\overline{Y}_{D_{i}=0,P_{t}=0}.\right)\end{split}$ (25)
To gain a more intuitive understanding of why they are equal, it is helpful to delve into the interpretation of the coefficients in the UN-DID regressions. Specifically, $\widehat{\lambda_{2}^{T}}$ represents the mean outcome of the treated group in the post-intervention period, while $\widehat{\lambda_{1}^{T}}$ signifies the mean outcome of the treated group in the pre-intervention period. Similarly, $\widehat{\lambda_{2}^{C}}$ corresponds to the mean outcome of the control group in the post-intervention period and $\widehat{\lambda_{1}^{C}}$ denotes the mean outcome of the control group in the pre-intervention period.
Proof: Rewriting Equation (6) and taking expectation conditional on $pre_{t}^{T}=1$ on both sides, we get
$Y^{T}_{i,t}=\lambda_{1}^{T}pre_{t}^{T}+\lambda_{2}^{T}post_{t}^{T}+\nu^{T}_{i,t}$
$\Rightarrow E[Y^{T}_{i,t}|pre_{t}^{T}=1]=E[\lambda_{1}^{T}pre_{t}^{T}|pre_{t}^{T}=1]+E[\lambda_{2}^{T}post_{t}^{T}|pre_{t}^{T}=1]+E[\nu^{T}_{i,t}pre_{t}^{T}=1]$
$\Rightarrow E[Y^{T}_{i,t}|pre_{t}^{T}=1]=\lambda_{1}^{T}E[pre_{t}^{T}|pre_{t}^{T}=1]+\lambda_{2}^{T}E[post_{t}^{T}|pre_{t}^{T}=1].$
Here, $E[\nu^{T}_{i,t}]=0$ by the strong exogeneity condition due to i.i.d sampling assumption. Because $pre_{t}^{T}=1$ is a dummy variable, we know that $E[pre_{t}^{T}|pre_{t}^{T}=1]=1$. By definition, we also know that $E[post_{t}^{T}|pre_{t}^{T}=1]=0$ because $post_{t}^{T}=1-pre_{t}^{T}$. Plugging in these values, we can rewrite the above equation as
$E[Y^{T}_{i,t}|pre_{t}^{T}=1]=\lambda_{1}^{T}.$
Therefore, the estimate of $\lambda_{1}^{T}$ is the mean outcome of the treated group in the pre-intervention period. Taking the expectation conditional on $post_{t}^{T}=1$ on both sides, we can also show that $E[Y^{T}_{i,t}|post_{t}^{T}=1]=\lambda_{2}^{T}$. The same can be shown for $\lambda_{1}^{C}$ and $\lambda_{2}^{T}$. Based on this result, we can say that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ is an estimate of the ATT shown in Equation (3). We also establish that $\widehat{\beta}_{3}=\left(\overline{Y}_{D_{i}=1,P_{t}=1}-\overline{Y}_{D_{i}=1,P_{t}=0}\right)-\left(\overline{Y}_{D_{i}=0,P_{t}=1}-\overline{Y}_{D_{i}=0,P_{t}=0}\right)$ from Equation (10). Here, $\overline{Y}_{D_{i}=k,P_{t}=l}$ represents the mean outcome of group $k$ in period $l$. When the law of large numbers holds, $\widehat{\beta}_{3}$ serves as the estimate of the ATT, as indicated in Equation (3). Assuming the sample is identical for both the conventional and UN-DID methods, both estimators should yield equal estimates of the ATT and their associated standard errors.
3.2 Monte Carlo Design
We conduct a set of Monte Carlo experiments with synthetic data to demonstrate the unbiasedness of the UN-DID estimator and its relative performance compared with the conventional DID estimator. The overall experiment is depicted in Figure 1, where we apply both conventional and UN-DID regression methods to the same simulated dataset to estimate ATTs and assess their properties. Note that, with poolable data, we can apply both the conventional and UN-DID methods to estimate the ATT. With unpoolable data, we are not able to run the conventional regression. Because these are synthetic data, they are poolable. Therefore, we can run both methods on the dataset.
Our generated dataset is intended to replicate samples from census data for Ontario and Quebec, specifically concentrating on individuals aged 65 and above. To be more precise, we aligned the age and gender distributions to closely resemble those observed in the populations of Quebec and Ontario in 2001 for individuals aged 65 years and older. We keep the gender constant for all individuals and age them by one year for the next nine years, resulting in our final dataset. Our outcome is a continuous variable. In our initial simulation, we created groups of 50 individuals each in both Ontario and Quebec, covering the period from 2000 to 2009. This results in a total of 500 observations evenly distributed over 10 years. We chose a small sample size to ensure the robustness of the results even when dealing with small datasets. For the simulation, we assume that Quebec is treated by some arbitrary policy in the year 2005. This approach ensures a balanced and symmetrical dataset, maintaining an equal number of observations in both pre- and post-treatment periods and in the treatment and control groups. We then repeat the simulations with roughly two thirds of the total observations in Ontario and one third of the total observations in Quebec. So, in a DGP with 500 observations, roughly 333 of the observations are in Ontario and 167 are in Quebec.
Figure 1: Schematic for Simulations
We also consider two data-generating processes to assess whether the estimates of the ATT align between conventional and UN-DID regressions. In the first case, we set the true treatment effect for the treated group at 0. In the second case, the true treatment effect for the treated group is set at 0.1. After generating the data, we perform both conventional and UN-DID regressions to estimate the ATT and access their unbiasedness properties. This is then repeated 1,000 times for each case. The data-generating process for the two cases are as follows:
  • Case 1: No treatment effect, no covariates:
    $Y^{g}_{it}=e^{g}_{it}.$
  • Case 2: Treatment effect, no covariates:
    $Y^{g}_{it}=0.1*DID^{g}_{it}+e^{g}_{it}.$
In all the simulations in the paper, the error term is idiosyncratic and follows a normal distribution, denoted as $e\sim\mathcal{N}(0,1)$. The $DID^{g}_{it}$ represents the interaction term between $P_{t}$ and $D_{i}$, corresponding to the post and treat dummy variables. The coefficient associated with this interaction term signifies the “true” effect of the treatment in both cases.
In this simulation study, the generated dataset is a balanced panel. The results displayed in the ensuring sections also apply to repeated cross sectional data. The results for time invariant covariate with repeated cross sectional data are comparable to the results with time varying covariates with panel data, as the distribution of the covariate varies over time due to different individualsbeing present in the data in each year.
3.3 Results: No Covariates
Figure 2: No Covariates with Unequal Sample Sizes: Kernel Density
The simulation study demonstrates that the UN-DID estimator is unbiased. To assess unbiasedness, we generate kernel density plots for both conventional DID and UN-DID estimates of the ATT on a shared axis. The kernel densities for unequal sample sizes are shown in Figure 2. In Panel (a), the kernel density estimates for Case 1 are presented, and in Panel (b), the kernel density estimates for Case 2 are presented. Notably, for Case 1, both kernels are centered around 0, aligning with the true value of the ATT. Similarly, for Case 2, both kernels are centered around 0.1, which is the true value of the ATT for Case 2. The blue line representing the conventional ATT lies exactly beneath the red line representing the UN-DID ATT (except at the tails) and is therefore not visible. This demonstrates that the ATTs of the two methods are exactly numerically equivalent. This finding demonstrates that both conventional and UN-DID estimation methods exhibit unbiasedness, as evidenced by the distributions of estimates being centered around the true values of the ATT. A similar pattern is found with equal sample sizes, in a figure not included.
In order to assess whether the UN-DID approaches the true value as sample size increases, we repeat the Monte Carlo simulation by increasing the sample size to 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes between both silos. Subsequently, we compute the mean squared error (MSE) between the estimated ATT using UN-DID and the true ATT from the underlying DGP according to Equation (26):
$MSE(\widehat{ATT})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{ATT_{j}}^{UN-DID}-ATT_{0}\biggr{)}^{2}.$ (26)
To determine the MSE of the standard errors, we initially calculate the true standard error of the ATT using the following formula:
$SE_{0}=\frac{\sum_{i}e^{2}}{(n-f)(DID-\overline{DID})^{2}}.$ (27)
Here, $\overline{DID}$ is the mean of the DID term from the generated dataset and $f$ is the degrees of freedom. In the case with no covariates, the degrees of freedom is set at 4, following the conventional regression shown in Equation (1). The MSEs for the standard errors are then computed using the formula shown in Equation (28):
$MSE(\widehat{SE})=\frac{1}{1000}\sum_{j}^{1000}\biggl{(}\widehat{SE}(\widehat{ATT_{j}})^{UN-DID}-SE_{0}\biggr{)}^{2}.$ (28)
The findings are shown in Figure 3, where the y-axis represents the mean squared error (MSE) and the x-axis corresponds to the sample size. The graph illustrates a decreasing trend in MSE for both the estimated ATT and the standard error as the sample size increases. The results for unequal sample sizes are illustrated in Figure 4. We observe that the MSE for the case with unequal sample sizes is higher than for the case with equal sample sizes but still converges to 0 as the sample size increases.
Figure 3: No Covariates with Equal Sample Sizes: Mean Squared Errors of UN-DID vs. true values
Figure 4: No Covariates with Unequal Sample Sizes: Mean Squared Errors of UN-DID vs. true values
4 UN-DID When Strong Parallel Trends Do Not Hold
The previous section was based on the assumption of strong parallel trends between the treated and the control silos. We now relax Assumption 2 and introduce the conditional parallel trends assumption, so that parallel trends is plausible only conditional on covariates.
Assumption 7: Conditional Parallel Trends
$\begin{split}\biggr{[}E[Y_{i}(0)|D_{i}=1,P_{t}=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=1,P_{t}=0,X_{i,t}]\biggr{]}\\ =\biggr{[}E[Y_{i}(0)|D_{i}=0,P_{t}=1,X_{i,t}]-E[Y_{i}(0)|D_{i}=0,P_{t}=0,X_{i,t}]\biggr{]}.\end{split}$
The conditional parallel trends assumption states that the evolution of outcomes between the treated and control groups are the same in the pre-intervention period, conditional on covariates. Under the conditional parallel trends assumption, we can also express the ATT, shown in Equation (3), as the ATT conditional on covariates:
$\begin{split}ATT(x)=&\biggl{[}E[Y_{i}|D_{i}=1,P_{t}=1,X_{i,t}=x]-E[Y_{i}|D_{i}=1,P_{t}=0,X_{i,t}=x]\biggr{]}\\ &-\biggl{[}E[Y_{i}|D_{i}=0,P_{t}=1,X_{i,t}=x]-E[Y_{i}|D_{i}=0,P_{t}=0,X_{i,t}=x]\biggr{]}.\end{split}$ (29)
With covariates, we can extend Assumption 6 such that matrices of the covariates, $X_{T}$ and $X_{C}$ for the treated and the control groups, respectively, can be stacked together into a common matrix $X$:
$Y_{T}=\begin{bmatrix}y_{T1}\\ y_{T2}\\ \vdots\\ y_{TN_{T}}\end{bmatrix},Y_{C}=\begin{bmatrix}y_{C1}\\ y_{C2}\\ \vdots\\ y_{CN_{C}}\end{bmatrix},X_{T}=\begin{bmatrix}x_{T1}\\ x_{T2}\\ \vdots\\ x_{TN_{T}}\end{bmatrix},X_{C}=\begin{bmatrix}x_{C1}\\ x_{C2}\\ \vdots\\ x_{CN_{C}}\end{bmatrix}$
$Y=\begin{bmatrix}Y_{T}\\ Y_{C}\end{bmatrix}=\begin{bmatrix}y_{T1}\\ y_{T2}\\ \vdots\\ y_{TN_{T}}\\ y_{C1}\\ y_{C2}\\ \vdots\\ y_{CN_{C}}\end{bmatrix},X=\begin{bmatrix}X_{T}\\ X_{C}\end{bmatrix}=\begin{bmatrix}x_{T1}\\ x_{T2}\\ \vdots\\ x_{TN_{T}}\\ x_{C1}\\ x_{C2}\\ \vdots\\ x_{CN_{C}}\end{bmatrix}$
Of course this stacking is not feasible when data are siloed. With unpoolable data, when we gain access to the data from the treated silo, we do not have any information on the data from the control group, or both $Y_{C}$ and $X_{C}$. Similarly, when we gain access to the data from the untreated silo, we do not have any information on the data for the treated silo, or both $Y_{T}$ and $X_{T}$. Because we do not observe the covariates from other groups, we are unable to use non-parametric methods such as the inverse probability weighting (IPW), outcome regression (OR), and the doubly robust methods (DR–DID) to estimate the ATT (Abadie, 2005; Sant’Anna and Zhao, 2020; Heckman et al., 1997).
Researchers typically incorporate covariates in conventional DID to ensure the plausibility of parallel trends given the covariates, particularly when strong parallel trends do not hold (Heckman et al., 1997; Abadie, 2005). Likewise, we are including only covariates that are required for conditional parallel trends to be plausible. For valid estimation of the ATT with covariates, we require that the covariates are correlated with the outcome but not with the treatment on the basis of the conditional independence assumption.
Assumption 8: Conditional Independence Assumption
[…]
The conditional independence assumption states that the treatment assignment is independent of potential outcomes after conditioning on a set of observed covariates (Masten and Poirier, 2018).
The literature on DID with covariates has largely focused on scenarios where covariates are time-invariant (Abadie, 2005; Callaway and Sant’Anna, 2021; Caetano et al., 2022; Roth et al., 2022; Sant’Anna and Zhao, 2020). When dealing with time-varying covariates, researchers frequently select a specific value of these covariates during a pre-treatment period, as if it were a time-invariant covariate (Caetano et al., 2022).
We will begin by examining the simplest case, where we introduce a single covariate into both the conventional DID and UN-DID regressions. With a single covariate, the conventional DID regression is expressed as
$Y_{i,t}=\beta_{0}+\beta_{1}D_{i}+\beta_{2}P_{t}+\beta_{3}P_{t}*D_{i}+\beta_{4}X_{i,t}+\epsilon_{i,t}.$ (31)
Likewise, we adjust the UN-DID regression to incorporate covariates:
[…]
Including time-invariant covariates does not complicate DID analysis, as demonstrated in the following subsection. However, with time-varying covariates, Caetano et al. (2022) has demonstrated that the two-way fixed effects (TWFE) DID regression can provide biased estimate of the ATT due to negative weighting and weight-reversal issues. In this paper, we recognize the complexity introduced by the heterogeneity in data distribution between silos, which can present challenges when analyzing siloed data (Li et al., 2022).
4.1 Common Causal Covariates
One important distinction between the conventional DID regression using poolable data in Equation (31) and the UN-DID regresssions using siloed data in Equations (32) and (33) is how the covariates are handled. In the conventional DID regression there is a single coefficient, $\beta_{4}$, whereas in the UN-DID regressions there are two $\lambda_{3}^{T}$ and $\lambda_{3}^{C}$. Whether or not this matters depends on the relationship between the covariate(s) and the outcome variable and if that varies across silos. Work in progress by Karim and Webb explores this issue in detail. That paper introduces the assumption of common causal covariates, which implies that the relationship between the covariate(s) and the outcome is the same in each silo. The simulations in this paper maintain that assumption. When that assumption is violated, the conventional DID estimator and the estimator from Callaway and Sant’Anna (2021) can be quite biased. Conversely, the UN-DID estimator, or the estimator using pooled data but with silo specific coefficients for covariates introduced by Karim and Webb is unbiased in these settings.
4.2 Equivalence of the Estimated ATT Between Conventional DID and UN-DID Methods with Covariates
Conventional DID Regression
For the conventional DID regression, we assume that Assumption 1 and Assumptions 3 to 7 hold. With a single covariate, the conventional regression is shown in Equation (31). In Equation (31), $\hat{\beta_{3}}$ is the parameter of interest, which is the estimate of the ATT conditional on covariates, and is the sample analogue of Equation (29). According to the FWL theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (35). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from Equation (34):
$P_{t}*D_{i}=\alpha_{0}+\alpha_{1}P_{t}+\alpha_{2}D_{i}+\alpha_{3}X_{i,t}+u_{i,t},$ (34)
$Y_{i,t}=\widehat{\beta_{3}}(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})+u^{\prime}_{i,t}.$ (35)
From Equation (34), we calculate residuals ($P_{t}*D_{i}-\widehat{P_{t}*D_{i}}$) for each observation based on their treatment status and time period. For treated observations in the pre-intervention period, the residual is given by $-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}$, while for the post-intervention period, it is $1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}$. Similarly, untreated observations have residuals of $-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}$ in the pre-intervention period and $-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}$ in the post-intervention period. Subsequently, $\widehat{\beta}_{3}$ from Equation (35) can be estimated using the following OLS formula:
$\begin{split}\widehat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)}{\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)^{2}}\end{split}$ (36)
The expression from Equation (36) is simplified by substituting the computed residuals into the numerator, as shown below:
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)+\\ &\sum_{i\in t=0,C}Y_{i,1}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}\right).\end{split}$
Similarly, by substituting the residuals and simplifying, the denominator of Equation (36) can be represented in the following manner:
$\begin{split}G=&\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,0}\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}X_{i,1}\right)^{2}.\end{split}$
Combining the numerator and the denominator, $\widehat{\beta}_{3}$ can be written as:
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}X_{i,1}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}X_{i,0}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}X_{i,1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}X_{i,0}\right)}{G}.\end{split}$ (37)
Now, let us interpret the meaning of $\widehat{\beta}_{3}$ in Equation (37). To simplify, consider the case where $X_{i,t}$ is a discrete variable and assign it a specific value, say $x$. When $x=0$, Equation (37) reduces to the same form as Equation (14), where $\widehat{\beta_{3}}$ represents the “difference-in-difference” of four unconditional means. When $x\neq 0$, Equation (37) can be expressed as follows:
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)}{G^{\prime}},\end{split}$ (38)
where
$\begin{split}G^{\prime}=&\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)^{2}\\ &+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{3}x\right)^{2}.\end{split}$
Continuing with the reasoning from the proof outlined in Appendix A, we can show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent to the reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, with $X_{i,t}=x$. In simpler terms, these represent the probabilities of the observation being in the treated group in their respective time periods, conditional on $X_{i,t}$. It is important to note that this equivalence holds with a sufficiently large sample size.
With smaller sample sizes, where the law of large numbers does not apply, therefore, $\widehat{\alpha}_{3}\neq\alpha^{0}_{3}$, the true effect of $X_{i,t}$ on the interaction term. Similarly, $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ represent the probabilities for the observations being in the control group in the post- and pre-intervention periods, respectively, conditional on $X_{i,t}$. As a result, $\hat{\beta_{3}}$ can be expressed as
$\begin{split}&\sum_{i\in t=1,T}P(D_{i}=1,P_{t}=1|X^{T}_{i,1}=x)Y^{T}_{i,1}-\sum_{i\in t=0,T}P(D_{i}=1,P_{t}=0|X^{T}_{i,0}=x)Y^{T}_{i,0}\\ &-\sum_{i\in t=1,C}P(D_{i}=0,P_{t}=1|X^{C}_{i,1}=x)Y^{C}_{i,1}+\sum_{i\in t=0,C}P(D_{i}=0,P_{t}=0|X^{C}_{i,0}=x)Y^{C}_{i,0}.\end{split}$ (39)
Assuming the law of large numbers holds because of large sample size, Equation (39) can be rewritten using the population analogue for means:
$\begin{split}\widehat{\beta_{3}}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}=x]-E[Y^{T}_{i,0}|X^{T}_{i,0}=x]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}=x]-E[Y^{C}_{i,0}|X^{C}_{i,0}=x]\right].\end{split}$ (40)
This equivalence does not hold when we have small sample sizes. In Equation (40), $\widehat{\beta}_{3}$ is the ATT conditional on covariates shown in Equation (29). If $X_{i,t}$ is a continuous random variable, Equation (39) can be rewritten as
[…]
This is also the ATT conditional on covariates and can be simplified to Equation (40) with large enough sample size.
UN-DID Regression
When conditional parallel trends is plausible, we can extend the UN-DID regressions shown in equations (6) and (7) to include covariates:
$\;Y^{j}_{i,t}=\lambda_{1}^{j}pre_{t}^{j}+\lambda_{2}^{j}post_{t}^{j}+\lambda^{j}_{3}X^{j}_{i,t}+\nu^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;j=\{T,C\},$ (42)
$\equiv\;Y^{j}_{i,t}=\gamma_{0}^{j}+\gamma_{1}^{j}post_{t}^{j}+\gamma_{2}^{i}X^{j}_{i,t}+\nu^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;j=\{T,C\}.$ (43)
Following the proof shown in Section 3.1, we know that $\widehat{\lambda}_{2}^{j}-\widehat{\lambda}_{1}^{j}$ from Equation (42) is equivalent to $\widehat{\gamma}_{1}^{j}$ from Equation (43). We proceed to derive the coefficient $\widehat{\gamma}_{1}^{j}$. Using the FWL theorem, it follows that $\widehat{\gamma}_{1}^{j}$ is equal to the coefficient of $(post^{j}_{t}-\widehat{post}^{j}_{t})$ in the regression presented in Equation (44). Here, $(post^{j}_{t}-\widehat{post}^{j}_{t})$ are the residuals from the model shown in Equation (45):
$post_{t}^{j}=\eta_{0}^{j}+\eta_{1}^{j}X^{j}_{i,t}+\omega^{j}_{i,t},\;\;\;\;\;\;\mbox{where}\;\eta_{0}^{j}=\overline{post}_{t}^{j},$ (44)
$Y_{i,t}^{j}=\gamma_{1}^{j}(post^{j}_{t}-\widehat{post^{j}_{t}})+\nu^{j^{\prime}}_{i,t}.$ (45)
Similar to the previous section, we compute the residuals $(post_{t}^{j}-\widehat{post}_{t}^{j})$ for each observation using Equation (45). For an observation in group $j$ during the pre-intervention period, the residual is given by $-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}$. Similarly, for an observation in group $j$ during the post-intervention period, the residual is expressed as $1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}$. Following this, we move on to derive $\widehat{\gamma}_{1}^{j}$ from Equation (44) using the OLS formula:
$\begin{split}\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}}\sum_{i\in t=0}Y_{i,0}\end{split}$
$\Rightarrow\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{F}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}}{F}\sum_{i\in t=0}Y_{i,0}.$ (46)
Here,
$F=\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}-\widehat{\eta}_{1}^{j}X^{j}_{i,t}\right)^{2}.$
Following the proof in Appendix A, we can show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}-\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}+\widehat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, $X_{i,t}=x$ (provided the sample size is large enough). Assuming that the law of large numbers hold, $\widehat{\gamma}^{j}_{1}$ in Equation (46) is the sample analogue of the following:
$\widehat{\gamma}_{1}^{j}=\left[E[Y^{j}_{i,1}|X^{j}_{i,t}]-E[Y^{j}_{i,0}|X^{j}_{i,0}]\right].$ (47)
From Equation (47), we can take the difference between $\widehat{\gamma}^{T}_{1}$ and $\widehat{\gamma}^{C}_{1}$ from the treated and untreated regressions, respectively. Equation (48) shows that the difference is the sample analogue of the ATT shown in Equation (29):
$\begin{split}\widehat{\gamma}_{1}^{T}-\widehat{\gamma}_{1}^{C}=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}]-E[Y^{T}_{i,0}|X^{T}_{i,0}]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}]-E[Y^{C}_{i,0}|X^{C}_{i,0}]\right]\end{split}$
$\begin{split}\Rightarrow(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})=\left[E[Y^{T}_{i,1}|X^{T}_{i,t}]-E[Y^{T}_{i,0}|X^{T}_{i,0}]\right]-\left[E[Y^{C}_{i,1}|X^{C}_{i,t}]-E[Y^{C}_{i,0}|X^{C}_{i,0}]\right].\end{split}$ (48)
It becomes evident when comparing UN-DID and conventional DID that both methods are capable of estimating the ATT with common slope parameters of the covariates across groups. Now, we will demonstrate that the ATT estimates derived from conventional and UN-DID regressions are equal with time-invariant covariates.
In the context of time-invariant covariates, the regression coefficients for $X_{i}$ in equations (34) and (44) are observed to be 0. This is a consequence of the conditional independence assumption, indicating that $X_{i}$ is independent of $D_{i}$ ($X_{i}\perp D_{i}$). Additionally, because $X_{i}$ is a time-invariant covariate, it is independent of $P_{t}$ as well ($X_{i}\perp P_{t}$). As a result, both $\widehat{\alpha}_{3}$ and $\widehat{\eta}^{j}_{1}$ are equal to 0. Therefore, $\widehat{\alpha_{3}}=\widehat{\eta}^{j}_{1}=0$. From this result, $\hat{\beta_{3}}$ in Equation (38) can be written as
$\begin{split}\widehat{\beta}_{3}=&\frac{\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)}{G}-\frac{\sum_{i\in t=0,T}Y_{i,0}^{T}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{G}\\ &-\frac{\sum_{i\in t=1,C}Y_{i,1}^{C}\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}\right)}{G}+\frac{\sum_{i\in t=0,C}Y_{i,0}^{C}\left(-\widehat{\alpha}_{0}\right)}{G}.\end{split}$ (49)
This is the same as the ATT in the absence of covariates for the conventional DID regression, shown in Equation (14). For the UN-DID regression, $\gamma_{1}^{j}$ in Equation (46) can be rewritten as
$\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.$ (50)
This is the same as the ATT in the absence of covariates for the UN-DID regression, shown in Equation (22). Therefore, the ATT estimates from both methods are the same.
The exact numerical equivalence is not maintained when time-varying covariates are present. As noted earlier, both $\widehat{\alpha}_{3}$ in Equation (34) and $\widehat{\eta}_{1}^{j}$ from Equation (44) are not equal to the true effect of $X_{i,t}$ because the law of large numbers does not hold with small sample size. However, as sample sizes increase, the estimates converge due to the law of large numbers. Thus, the equivalence between conventional DID and UN-DID estimates with time-varying covariates holds primarily with large sample sizes.
4.3 Monte Carlo Design with Time-Invariant Covariates
In this section, we continue to use the same synthetic dataset described in Section 3.2. We introduce a modification to our data-generating process, where the outcome is now influenced by a time-invariant covariate, denoted by $X_{i}$. In our simulations, the time-invariant covariate is represented by a binary variable indicating gender. Specifically, it assumes a value of 1 if the individual is female, and 0 otherwise. Similar to the previous section, we present two cases. In the first case, we set the true treatment effect at 0. In the second case, we set the true treatment effect at 0.1. Additionally, we assign a common slope parameter of 0.5 to the covariate in both the treated and the control silos.
After generating the data, we estimate the ATT using both the conventional and UN-DID regressions to evaluate their properties with time-invariant covariates. This is repeated 1,000 times for each case, following the methodology employed above. The data-generating processes for the two cases are as follows:
  • Case 3: No treatment effect, time-invariant covariate
    $Y^{g}_{it}=0.5X^{g}_{i}+e^{g}_{it}.$
  • Case 4: Treatment effect, time-invariant covariates
    $Y^{g}_{it}=0.1*DID^{g}_{it}+0.5X^{g}_{i}+e^{g}_{it}.$
4.4 Results: Time-Invariant Covariate
In subsection 4.2, we showed that both the conventional DID and the UN-DID methods can recover numerically equivalent values of the ATT when implemented on a poolable dataset. However, the proofs are limited to the estimation of the ATT and do not show the equivalence of the standard errors between the two methods. Through the Monte Carlo Simulation study, we assess the equivalence of the estimated standard errors between the conventional and the UN-DID methods using a scatter plot, shown in Figure 5. Panels (a) and (b) depict the relationships between the conventional DID standard error (y-axis) and the UN-DID standard error (x-axis) for Case 3 and Case 4 with unequal sample sizes, respectively. Reviewing the scatter plots, we note that the standard errors lie along the $45^{\circ}$ line but do not align perfectly with it. This indicates that the standard errors are equivalent but not numerically equal. The same results hold for equal sample sizes (figure not shown).
Figure 5: Time-Invariant Covariate with Unequal Sample Sizes: Scatter Plots
We assess the unbiasedness of the UN-DID estimator with unequal sample sizes and time-invariant covariates using kernel densities, illustrated in Figure 6. The kernel densities show that the distribution of ATTs is centered on the true ATT value, implying unbiasedness with time-invariant covariates. As in Figure 2, the blue line is not visible as it is perfectly underneath the red line (except at the tails), demonstrating numerical equivalence.
Figure 6: Time-Invariant Covariates with Unequal Sample Sizes: Kernel Density
We hypothesize that the standard errors are not numerically equivalent due to small sample sizes. To verify that the two methods converge to the true standard error as sample size increases, we repeat the simulation with sample sizes of 1,000, 2,000, 4,000, 8,000, 10,000, and 50,000 while maintaining equal sample sizes in both silos. We then compute the MSE for both the ATTs and their associated standard errors for each pair, looking for evidence of convergence in the results shown in Figure 7. The MSEs on the y-axis are plotted against the sample sizes on the x-axis. The MSEs between the UN-DID and the true ATT are illustrated by the red lines in Figure 7. On the same axis, we plot the MSEs comparing the conventional DID to the true ATT (the blue lines in Figure 7) and the MSEs between the conventional DID and UN-DID methods (the green lines in Figure 7).
Figure 7 depicts the MSE of ATTs for Case 3 in Panel (a) and for Case 4 in Panel (b). Panels (c) and (d) showcase the MSE of standard errors for Case 3 and Case 4, respectively. We observe that both ATT and standard error MSEs decrease with larger sample sizes, confirming that both the conventional and UN-DID methods converge to the true values of both the ATT and the standard errors as sample size increases. The blue lines (which illustrate the MSEs for the conventional DID vs. true values of ATT and SE) are not visible, as they lie perfectly underneath the red line (MSE for the UN-DID vs. true values of ATT and SE). This also shows the equivalence between the conventional DID and UN-DID ATT and SEs with time-invariant covariates. Figure 7 also illustrates a near zero MSE between the conventional DID and UN-DID methods, indicating the equivalence between the two estimators. The same pattern is observed in a case with unequal sample sizes, as shown in Figure 8.
Figure 7: Time-Invariant Covariate with Equal Sample Sizes: Mean Squared Errors
Figure 8: Time-Invariant Covariate with Unequal Sample Sizes: Mean Squared Errors
4.5 Monte Carlo Design with Time-Varying Covariates
In this section, we assess the performance of the UN-DID estimator with time-varying covariates. The Monte Carlo simulation design closely resembles the simulation of the previous subsection, however we replace the time-invariant covariate with a time-varying covariate, denoted by $X_{it}$. In our simulations, the time-varying covariate is age, which is unaffected by treatment even though it changes with time. Similar to the previous subsection, we analyze two cases where the true effect of the treatment is 0 and 0.1, respectively. A common slope parameter of 0.5 is assigned to the covariate in both the treated and the control silos. The DGPs for the two cases are as follows:
  • Case 5: No treatment effect, time-varying covariate
    $Y^{g}_{it}=0.5X^{g}_{it}+e^{g}_{it}.$
  • Case 6: Treatment effect, time-varying covariate
    $Y^{g}_{it}=0.1*DID^{g}_{it}+0.5X^{g}_{it}+e^{g}_{it}.$
4.6 Results: Time-Varying Covariate
In subsection 4.2, we showed that both the conventional and UN-DID methods effectively identify the ATT in the presence of time-varying covariates with common slope parameters. The ATT estimates are close (not exactly numerically equal), and converge to the true value in larger samples due to the law of large numbers. Therefore, the equivalence between the conventional DID and the UN-DID estimates holds in larger samples.
To assess the equivalence of ATT estimates and their associated standard errors, we created scatter plots (Figure 9). In Panel (a), we compare the conventional DID ATT (y-axis) with the UN-DID ATT (x-axis) for Case 5 with unequal sample sizes. Panel (b) shows a similar comparison for Case 6. Additionally, panels (c) and (d) illustrate the relationship between the conventional DID standard errors (y-axis) and the UN-DID standard errors (x-axis) for cases 5 and 6, respectively. Upon inspecting the scatter plots, we observe that the points representing ATT estimates and their associated standard errors align along the $45^{\circ}$ line. However, they are not perfectly aligned, suggesting the ATT estimates are not numerically equal but are equivalent. This pattern is consistent across both cases. We also find a similar pattern with equal sample sizes (figures not shown).
Figure 9: Time-Varying Covariate with Unequal Sample Sizes: Scatter Plots
We assess the unbiasedness of the UN-DID estimator with unequal sample sizes and time-varying covariates using kernel densities, illustrated in Figure 10. The kernel densities reveal that the distribution of ATTs is centered around the true ATT value, indicating UN-DID’s unbiasedness with time-varying covariates. Compared with Figures 2 and 6, there are now some small differences between the conventional DID and UN-DID ATTs.
In Figure 11, we observe convergence toward zero in the MSEs for both the conventional DID and UN-DID methods compared to the true value as sample size increases with equal sample sizes. The same patterns are observed for unequal sample sizes between silos (Figure 12).
Figure 10: Time-Varying Covariates with Unequal Sample Sizes: Kernel Density
Figure 11: Time-Varying Covariate with Equal Sample Sizes: Mean Squared Errors
Figure 12: Time-Varying Covariate with Unequal Sample Sizes: Mean Squared Errors
4.7 Differing Covariates by Silo
One interesting feature of siloed data is that the information may be collected and coded differently according to the processes in place in each jurisdiction. Each silo is primarily interested in analyzing the data in isolation for their own purposes, not in comparisons with other jurisdictions. It is therefore likely that not all variables will be standardized or harmonized across silos. UN-DID is designed for settings where data are unpoolable in the sense that they cannot be combined on one server or in one regression, but where the information they capture is similar enough that fundamental issues of data harmonization are not a problem. It is convenient, therefore, that UN-DID can accomodate settings where data are sufficiently harmonized so that variables capture the same concepts across silos, even if they are coded somewhat differently.
Consider, for instance, a variable like family income. Ideally this variable includes the same types of income (wages, investment earnings, government transfers, etc.) and is continuous and not-censored in both silos. However, it is plausible that both silos record family income as categorical: one using bins in increments of \$10,000, while the other uses increments of \$25,000. This would cause issues for conventional DID, as the categories are non-overlapping across silos without a problematic degree of aggregation. With UN-DID, the silo-specific coefficients can handle the different categories easily, each first-difference can be estimated conditional on the covariate, and the conditional ATT and SEs can also be estimated. While data must be sufficiently harmonized to capture the same underlying concepts and the common causal covariate assumption must hold, UN-DID can more flexibly accomodate differences in datasets across data silos than conventional DID.
5 UN-DID with Staggered Adoption
In this section, we relax Assumption 5 to accommodate a scenario where multiple groups adopt treatment at different points in time, creating a staggered adoption framework. We also impose an additional restriction called absorptive state, which implies that, once a unit is treated, it remains treated for the rest of the period of the study. Similar to Callaway and Sant’Anna (2021), UN-DID with staggered adoption will decompose the dataset into several $2\times 2$ group–time cells. Staggered adoption means that there are at least three silos under analysis. Section 6 gives additional guidance for working with multiple silos.
We introduce some additional notation following Callaway and Sant’Anna (2021). Let $G_{i,g}=1$ if unit $i$ is treated at time $g$, and 0 otherwise. These will be referred to as the treatment start-time dummies. Following Callaway and Sant’Anna (2021), we group together each individual into $g\in G$ cohorts based on the treatment start-time dummies. For a particular group $g$, any period where $t<g$ are called pre-treatment periods. Under the no-anticipation assumption, the ATT(g,t)s for pre-treatment periods are 0.
Each group–time cell contains a group that is currently treated and a group that has not yet been treated. If treated and control observations within this cell are in separate data silos, we can implement UN-DID to estimate the ATT(g,t). The pre-intervention period in each $2\times 2$ cell is defined as the time period right before the treated group is treated ($g-1$). For each $2\times 2$ group–time cell, we will run the following two regressions:
[…]
After these two regressions are estimated, the ATT(g,t) is
$\widehat{ATT(g,t)}=(\zeta^{T}_{2}-\zeta^{T}_{1})-(\zeta^{C}_{2}-\zeta^{C}_{1}).$ (53)
The above process is repeated for all relevant $2\times 2$ group–time cells. These ATT(g,t) cells are then aggregated together to get an overall estimate of the ATT based on Equation (54). $w_{g,t}$ are the weights associated with each of the ATT(g,t) cells.
$\widehat{ATT}=\sum_{g=2}^{G}\sum_{t=2}^{\mathcal{T}}1\{g\leq t\}w_{g,t}ATT(g,t).$ (54)
Common weighting schemes include equal weighting for each cell and population-based weighting. Refer to Callaway and Sant’Anna (2021) and Xiong et al. (2023) for details.
6 Multiple Silos
Up until Section 5, we have concentrated on settings in which there are only two silos. However, even with common timing it is typically the case that many treated and many comparison silos will be used. In that case, it is easier to estimate the average treatment effect by means of a second stage regression. The details differ slightly depending on whether there is common or staggered adoption.
6.1 Common Adoption
Imagine that there are two treated silos, A and B, and one control silo C. For these purposes, consider that the various silos contain the individual level observations, and that the data sharing agreements permit summary statistics to be extracted, but no new information can be added to a silo’s secure environment. In that case, calculation of the treatment effect must occur on a fourth computer, call this a server. Let us also assume that we are not interested in estimating dynamic treatment effects, so that aggregating time points into pre- and post-intervention will suffice. In that case, we need three pieces of information from each silo: the difference between the pre-period and the post-period, as estimated in Equation (42), the standard error for this difference, and possibly the weight for the silo. We typically recommend the weight for silos $s$ of $w_{s}=N^{post}_{s}/(N^{post}_{s}+N^{pre}_{s})$.
The first stage of the analysis creates a dataset which contains this information, along with treatment status for each of the three silos. In this case, the dataset would have three observations, and four variables: $\textrm{diff}_{s}$ is the within silo pre-post difference, potentially after controlling for covariates; $se_{s}$ is the standard error for $\textrm{diff}_{s}$; $w_{s}$ is the silo weight; and $d_{s}$ is whether that silo is treated $d_{s}=1$ or not $d_{s}=0$. We can calculate the ATT by estimating the regression:
$\textrm{diff}_{s}=\alpha+\beta d_{s}+\epsilon_{s}.$ (55)
In this regression, $\beta$ is the ATT, which will give the same estimate that one would get from a regression using poolable data when the groups are of equal size. When the groups are of unequal size, a weighted least squares version of the regression is needed to match the regression with pooled data. This second stage regression approach is particularly useful for cluster robust inference. This is being explored in a companion paper in development by the authors of this paper.
6.2 Staggered Adoption
Staggered adoption means an additional computational burden with multiple silos. Each individual ATTg,t can be calculated as before, and then an overall ATT can be estimated as a weighted average (see Section 5). To fix ideas, consider the simplest staggered adoption setting, where there are three groups (A,B,C) and three time periods (1,2,3). Group A is first treated in period 2 ($g=2$). Group B is first treated in period 3 ($g=2$). Group C is never treated ($g=\infty$). In this 3$\times$3 example, without covariates, and using the never-treated cells as controls, the three $ATT_{gt}$s are:
$\begin{array}[]{rcl}ATT_{22}&=&(\bar{Y}_{A2}-\bar{Y}_{A1})-(\bar{Y}_{C2}-\bar{Y}_{C1})\\[8.0pt] ATT_{23}&=&(\bar{Y}_{A3}-\bar{Y}_{A1})-(\bar{Y}_{C3}-\bar{Y}_{C1})\\[8.0pt] ATT_{33}&=&(\bar{Y}_{B3}-\bar{Y}_{B2})-(\bar{Y}_{C3}-\bar{Y}_{C2})\end{array}$
In this example, we need to calculate differences in mean outcomes within each silo for different sets of years for each silo. Specifically, we need to estimate the difference between year 2 and 1, and 3 and 1 for silo A. For silo B, we only need the difference between year 3 and 2. Finally, for silo C, we need all three of these differences. Similar to the multiple silo with common timing example, we propose storing these differences, and their standard errors, in a separate matrix. The staggered treatment timing requires us to also keep track of the $g$ and $t$ for each difference that we calculate, as well as whether a given silo is a treatment or control silo. To fill this matrix, each of the within-silo, two-year differences can be calculated using a separate regression. Table 1 illustrates this matrix for a simple 3$\times$3 example.
Table 1: Second Stage Matrix
Contrast Silo d G T diff
A22 A 1 2 2 $\bar{Y}_{A2}-\bar{Y}_{A1}$
A23 A 1 2 3 $\bar{Y}_{A3}-\bar{Y}_{A1}$
B33 B 1 3 3 $\bar{Y}_{B3}-\bar{Y}_{B2}$
C22 C 0 2 2 $\bar{Y}_{C2}-\bar{Y}_{C1}$
C23 C 0 2 3 $\bar{Y}_{C3}-\bar{Y}_{C1}$
C33 C 0 3 3 $\bar{Y}_{C3}-\bar{Y}_{C2}$
The various ATTg,t terms, can then be estimated using this matrix estimating this regression:
$\textrm{diff}_{s,g,t}=\alpha+\beta d_{s,g,t}+\epsilon_{s,g,t}\textrm{ if }G=g\textrm{ and }T=t.$
In this regression $\beta$ estimates the ATTg,t given by the sample restriction. For the 3$\times$3 example, the three restrictions needed would be G=2, T=2; G=2, T=3; and G=3, T=3. We have developed a software package, UNDID which implements this algorithm. The package was built in Julia, but also has implementations in R and Stata, the package is available here: https://github.com/ebjamieson97/Undid.jl.
7 Empirical Examples
We now consider two empirical examples, one with common adoption timing and one with staggered adoption. We include these examples to highlight the properties of the UN-DID and conventional DID methods. In contrast to our Monte Carlo simulations, we do not know the true effects of the policies we analyze here. Using survey data that is pooled across US states, we can estimate the ATT using both conventional DID and UN-DID (treating the data as if it were siloed by state), and then compare the ATTs and SEs estimated using these two approaches. Our purpose here is to compare the two methods with real-world data - not to answer the substantive question. Accordingly, we diverge somewhat from the designs in the original analyses, and our analysis falls well short of complete empirical analyses of the substantive treatment effects.
7.1 The Impact of Medical Marijuana Legalization on Body Mass Index
We revisit the question asked in Sabia et al. (2017) of how access to medical marijuana impacts body mass index. The original study investigates the role of medical marijuana laws in 21 US states and finds that access to medical marijuana decreases the likelihood of being obsese. To assess the performance of UN-DID in a 2$\times$×T setting with common treatment timing, we explore the impact of the District of Columbia’s (DC) medical marijuana law, using New York State as the control.
Data come from the Behavioral Risk Factor Surveillance System (BRFSS). This repeated cross-sectional survey is pooled across US states but can also be analyzed as if the data were siloed by state. We use data from 2006–2013 from New York and DC, which enacted and enforced their medical marijuana law in 2010. We therefore have a $2\times 8$ design with an equal number of pre-treatment and post-treatment years. We restrict our attention to respondents aged 25 or older, resulting in a final sample of 84,783 observations.
With these data, we estimate treatment effects using four different models: a conventional DID regression and UN-DID, each with and without controls. The control variables are female and categorical variables for age, race, and education, which are time-varying covariates in a repeated cross-sectional dataset. The UN-DID estimates are calculated by artificially siloing the data by state. The conventional DID model is shown in Equation (56).
$\footnotesize\text{BMI}_{ist}=\beta_{0}+\beta_{1}\text{post}_{t}+\beta_{2}\text{treat}_{s}+\beta_{1}\text{did}_{st}+\delta_{1}\text{female}_{ist}+\delta_{2}\text{RACE}_{ist}+\delta_{3}\text{AGE}_{ist}+\delta_{4}\text{EDUC}_{ist}+\epsilon_{ist}.$ (56)
We also the UN-DID ATT by first estimating the following model separately for New York and DC.
$\footnotesize\text{BMI}_{ist}=\beta_{1}\text{post}_{t}+\beta_{2}\text{post}_{t}+\delta_{1}\text{female}_{ist}+\delta_{2}\text{RACE}_{ist}+\delta_{3}\text{AGE}_{ist}+\delta_{4}\text{EDUC}_{ist}+\epsilon_{ist}.$ (57)
Table 2 presents the ATT and standard error estimates. As expected, without covariates the conventional DID and UN-DID yield the same estimates: an estimated ATT of 0.2560 and a standard error of 0.0853. When we add covariates, the estimated coefficient drops to 0.1322 for the conventional DID and 0.1204 for UN-DID. These estimates are neither substantively nor statistically different from one another. The small difference between them is expected given that the models include 25 covariates which are restricted to common slopes with conventional DID and allowed to vary as silo-specific slopes with UN-DID. The standard errors are also slightly different, but substantively equivalent: 0.0807 for the conventional DID and 0.0811 for UN-DID.
This empirical example demonstrates the strong performance of UN-DID relative to conventional DID in a 2$\times$×T setting with common or single treatment timing. We reiterate that our effect estimates in this simplified example should not be interpreted as responding to the substantive question. However, this analysis using real-world, artificially siloed data demonstrates that UN-DID can accurately estimate the ATT and SE as well as conventional DID, both with and without covariates.
Table 2: Effect of DC’s Medical Marijuana Legalization on BMI
Conventional DID UN-DID Conventional DID UN-DID
Coefficient 0.2560 0.2560 0.1322 0.1204
Standard error 0.0853 0.0853 0.0807 0.0811
Covariates No No Yes Yes
7.2 The Impact of Merit Scholarships on College Attendance
To highlight the use of UN-DID in a staggered adoption setting, we revisit the analysis of merit scholarships from Conley and Taber (2011). These authors estimate the average impact of 10 state-level merit scholarship programs with different adoption dates. These scholarships vary in their design, generally awarding scholarships to high-achieving high school students to enroll in a college within their home state (see Deming and Dynarski (2010) and the references therein for details).
Data come from the Current Population Survey (CPS), another repeated cross-sectional survey that is pooled across US states. There are 51 “states” (including DC) and 12 years (1989-2000) in the dataset, for a total of 42,161 individual-level observations. We estimate the ATT of a state’s adoption of a merit scholarship program on the likelihood of a student in that state graduating college. We first estimate this effect individually for each of the 10 treated states, and then combine those for an aggregate ATT. This allows us to compare 11 distinct treatment effect estimates using both UN-DID and conventional DID.
Ten states adopt a merit scholarship across varying adoption dates. Four states adopt the scholarship in 1991, 1993, 1996, and 1999, respectively, while two states each adopt in 1997, 1998, and 2000. For each treated state, one control state was chosen from the set of 41 untreated states to roughly match the number of treated and untreated observations in each model. We use only one treated state to highlight the properties of the UN-DID estimator rather than to estimate the best possible counterfactuals. We estimate the treatment effects both with and without covariates. Specifically, using only data from treated state $T$ and control state $C$, we estimate the model:
$\text{college}_{ist}=\beta_{1}\text{post}_{t}+\beta_{2}\text{treat}_{s}+\beta_{1}\text{did}_{st}+\delta_{1}\text{black}_{ist}+\delta_{2}\text{asian}_{ist}+\delta_{3}\text{male}_{ist}+\epsilon_{ist}.$ (58)
The coefficients $\delta_{1},\delta_{2}$, and $\delta_{3}$ are estimated only when the covariates are included. The remaining variables are defined the standard way: $treat_{s}$ is set equal to 1 for observations in state $T$ and is set equal to 0 for observations in state $U$. Similarly, $post_{t}$ is set equal to 1 for years greater than and equal to the year when state $T$ first offered a merit scholarship. Finally, $did_{st}$ is the product of $treat_{s}$ and $post_{t}$. We estimate treatment effects, or $\widehat{did}_{st}$, and standard errors using the pooled dataset.
We then artificially silo the data to estimate treatment effects and standard errors using UN-DID. Specifically, for each treated silo we estimate
$\text{college}_{ist}=\beta_{1}\text{pre}_{st}+\beta_{2}\text{post}_{st}+\delta_{1}\text{black}_{ist}+\delta_{2}\text{asian}_{ist}+\delta_{3}\text{male}_{ist}+\epsilon_{ist}\text{ for }s\in(C,T).$ (59)
Here, $\text{pre}_{t}$ is equal to one in the years before the merit program was adopted, and equal to zero afterwards. Conversely, $\text{post}_{t}$ is equal to zero in the years before the merit program was adopted, and equal to one afterwards. We estimate the same regression for the control state used for each treated state. We then calculate the treatment effect using Equation (8) and the standard error using Equation (9).
The results are found in Table 3. Panel A shows the estimates for the models without covariates. The treatment effect and standard error estimates are identical to the fourth decimal place using conventional DID with the pooled data or UN-DID with the unpooled data. These results match the simulation results presented in Section 3.3.
Panel B of Table 3 shows the treatment effect estimates conditional on covariates. Here, some of the individual state treatment effect estimates are very similar using conventional DID and UN-DID. For instance, for state 88, the conventional DID effect estimate is 0.1811 and the UN-DID estimate is 0.1819. However, some of the estimates are not as close. The estimated treatment effect for state 61 is 0.0370 using conventional DID and 0.0314 with UN-DID. Even though the difference between the estimates is larger, they are not statistically different and may not be substantively different from each other. The estimates of the standard errors differ more between conventional DID and UN-DID. Again, these differences are not large enough to impact inference in this example. These findings are consistent with the results in Section 4.5.
Lastly, we estimate average treatment effects using data from all groups and all necessary years per group. We do this using two methods, both including covariates. We estimate the ATT using the CSDID command, which implements the Callaway and Sant’Anna (2021) procedure. We again pretend that each state’s data are siloed, and estimate the ATT using UN-DID. We do this using our UNDID command, which excludes forbidden comparisons. Simplifying the procedure, we estimate an ATT for each post treatment period for each treated state. This is done by estimating a model like in Equation (59) for each treated state, but using only one pre-treatment and one post-treatment year. Using those same years, we estimate the same model for each of the control states. The state-specific first differences $\beta_{2}-\beta_{1}$ are then calculated, and added to a table, like Table 1 as described in Section 5. We can then estimate each ATTg,t by running a properly specified regression like the one in Equation (55). For both CSDID and UN-DID we estimate one ATT by taking a simple average of the ATTg,t and one by averaging over the $g$ groups. Additionally, we calculate cluster-robust standard errors, assuming state level clustering, for all ATT estimates. The UN-DID standard errors use the jackknife discussed in Section 2.1.
The aggregate ATT estimates are shown in Panel C of Table 3. We find the simple average ATT estimates from the two procedures to be quite similar. The UN-DID ATT is 0.0485 and the CSDID ATT is 0.0464. The two standard errors are also very close to one another. The group-averaged ATTs are not as similar to one another, but still close. The UN-DID ATT is 0.0459, remaining close to the simple average ATT. The CSDID ATT drops to 0.0339. The standard errors are also not as similar, with the UNDID standard error equal to 0.0188 and the CSDID one equal to 0.0118. Given the UN-DID standard error is from a cluster jackknifed and the CSDID standard error is not, it makes sense that it is larger as the 51 states are unbalanced in the number of observations. This example, using 51 artificial silos, shows how accurately the UN-DID procedure can estimate an ATT.
Table 3: Estimates from Merit Example
Panel A: Without Covariates
Treated Control Year DID UN-DID DID SE UN-DID SE
71 73 1991 00-0- 0.0242 00-0- 0.0242 0.0755 0.0754
58 46 1993 0.1386 0.1386 0.0590 0.0590
64 54 1996 0.0182 0.0182 0.0619 0.0619
59 23 1997 00-0- 0.0073 00-0- 0.0073 0.0362 0.0362
85 86 1997 0.1318 0.1318 0.0601 0.0601
57 32 1998 00-0- 0.0291 00-0- 0.0291 0.0747 0.0747
72 55 1998 00-0- 0.0645 00-0- 0.0645 0.0658 0.0658
61 62 1999 0.0335 0.0335 0.0816 0.0816
34 33 2000 00-0- 0.0305 00-0- 0.0305 0.0666 0.0666
88 11 2000 0.1595 0.1595 0.1337 0.1337
Panel B: With Covariates
Treated Control Year DID UN-DID DID SE UN-DID SE
71 73 1991 00-0- 0.0141 00-0- 0.0137 0.0750 0.0871
58 46 1993 0.1501 0.1523 0.0582 0.0734
64 54 1996 0.0077 0.0028 0.0614 0.0820
59 23 1997 00-0- 0.0179 00-0- 0.0139 0.0360 0.0441
85 86 1997 0.1309 0.1277 0.0597 0.0726
57 32 1998 00-0- 0.0480 00-0- 0.0417 0.0733 0.0876
72 55 1998 00-0- 0.0704 00-0- 0.0685 0.0655 0.0793
61 62 1999 0.0370 0.0314 0.0806 0.0918
34 33 2000 00-0- 0.0282 00-0- 0.0303 0.0655 0.0709
88 11 2000 0.1811 0.1819 0.1335 0.1428
Panel C: Full Sample
Agg. UN-DID ATT UN-DID SE CSDID ATT CSDID SE
group 0.0459 0.0188 0.0339 0.0118
simple 0.0485 0.0110 0.0464 0.0117
8 Conclusion
Difference-in-differences is commonly used to estimate treatment effects, particularly for policies or other interventions that vary at the jurisdictional (e.g., state, country) level. However, this approach was previously infeasible in settings where the data are siloed by jurisdiction. In this paper, we propose the unpoolable DID (UN-DID) method, assess its performance, and demonstrate its utility in settings with siloed data. We also formalize the assumption of data poolability in the DID context, which was previously implied but not directly articulated. Specifically, we show that UN-DID is an unbiased and reliable extension to the conventional regression-based DID method in settings with unpoolable data.
We began by presenting analytical proofs to demonstrate the equivalence of UN-DID and conventional DID, both with and without covariates. We then extended our assessment to simulated panel data using a series of Monte Carlo simulations (incorporating a range of DGPs and sample sizes). These experiments permitted a direct assessment of UN-DID’s performance and unbiasedness compared to both conventional DID and the true treatment effects. To better understand UN-DID’s performance with real-world data, we concluded by comparing the performance of UN-DID and conventional DID using pooled data from existing empirical examples, treating these data as unpoolable.
Across all assessments, we showed that UN-DID and conventional DID consistently return identical treatment effect estimates and standard errors in models without covariates. The ATTs remain unbiased when we include time-invariant and time-varying covariates. The SEs are also unbiased but not numerically equal, converging to the true value as the sample size increases. These findings suggest that UN-DID is a valid approach to DID when data are unpoolable.
Our work contributes to a growing body of literature developing DID methods for use in increasingly complex policy and data environments. UN-DID offers an important methodological tool for researchers assessing multi-jurisdictional (or cross-jurisdictional) policy effects. By enabling the use of DID estimation in siloed data settings, UN-DID effectively opens the door to new research questions and new evidence on policy impacts.
The UN-DID software packages in Stata, R, and Julia will facilitate the practical application of UN-DID by a range of different end-users; the accompanying user guide will further support this aim while additionally showcasing the use of UN-DID in a truly siloed setting. Taken together, we expect our development of UN-DID (both as an analytical method and an associated software package) to catalyze innovative cross-jurisdictional policy evaluation in siloed data settings, thereby eliminating a common barrier for an otherwise widely used and powerful method.
A Interpretation of the Coefficients in Equation (11)
To begin, let us rewrite Equation (11) and proceed to derive the values for $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$:
$P_{t}*D_{i}=\alpha_{0}+\alpha_{1}P_{t}+\alpha_{2}D_{i}+u_{i,t}.$ (A1)
With a constant, we prove that the coefficient of a dummy variable represents the difference in means when the dummy variable takes on a value of 1 compared to when it takes on a value of 0.
General Proof: Let $M$ be a continuous outcome variable and $N$ be a binary dummy variable. The regression of $M$ on $N$ is expressed as
$M=\kappa_{0}+\kappa_{1}N+\upsilon.$ (A2)
By taking expectations, once when $N=1$ and once when $N=0$, we obtain
$E[M|N=1]=\kappa_{0}+\kappa_{1}$ (A3)
$\Rightarrow E[M|N=0]=\kappa_{0}.$ (A4)
With strict exogeneity, we know that $E[\upsilon|N=1]=0$ and $E[\upsilon|N=0]=0$. By subtracting Equation (A4) from Equation (A3), it becomes evident that
$\kappa_{1}=E[M|N=1]-E[M|N=0].$ (A5)
Based on this proof, we can write $\alpha_{1}$ as
$\alpha_{1}=E[P_{t}*D_{i}|D_{i}=1]-E[P_{t}*D_{i}|D_{i}=0].$ (A6)
When $D_{i}=0$, $P_{t}*D_{i}=0$. Therefore, $E[P_{t}*D_{i}|D_{i}=0]=0$. So we can rewrite Equation (A6) as
$\alpha_{1}=E[P_{t}*D_{i}|D_{i}=1]=E[P_{t}]=Pr(P_{t}).$ (A7)
Because $P_{t}$ is a dummy variable, $E[P_{t}]$ is the fraction of individuals in the post-intervention period. We can similarly show that $\alpha_{2}=E[D_{i}]=Pr(D_{i})$. To derive $\alpha_{0}$, let us take the unconditional expectation of Equation (A1):
$E(P_{t}*D_{i})=\alpha_{0}+\alpha_{1}\cdot E(D_{i})+\alpha_{2}\cdot E(P_{t})$ (A8)
$\Rightarrow Pr(P_{t})\cdot Pr(D_{i})=\alpha_{0}+Pr(P_{t})\cdot Pr(D_{i})+Pr(D_{i})\cdot Pr(P_{t}).$ (A9)
This follows from the fact that the expected value of a dummy variable is the fraction of the total sample with the dummy variable value = 1. The proof is shown below. We have also shown in Equation (A7) that $\alpha_{1}=Pr(D_{i})$ and $\alpha_{2}=Pr(P_{t})$.
Proof: The expected value of a binary dummy variable $D_{i}$ is given by
$E[D_{i}]=\frac{\sum_{i=1}^{n}D_{i}}{n}.$ (A10)
Because $D_{i}$ can take only the value of 0 or 1, we can express it as the count of observations, where $D_{i}=1$:
$E[D_{i}]=\frac{\sum_{i=1}^{n}D_{i}}{n}=\frac{\text{Number of observations where }D_{i}=1}{n}.$ (A11)
The same holds for $E(P_{t})$. We can now re-arrange and simplify Equation (A9) as follows:
$\alpha_{0}=-Pr(P_{t})\cdot Pr(D_{i}).$ (A12)
In other words, $\alpha_{0}$ is minus the fraction of individuals with both $P_{t}=1$ and $D_{i}=1$, $\alpha_{1}$ is the fraction of individuals in the post-intervention period, and $\alpha_{2}$ is the fraction of individuals in the pre-intervention period.
With equal sample sizes, $Pr(P_{t})=Pr(D_{i})=1/2$. This follows from the fact that approximately half of the total observations are in the post-intervention period and approximately half of the observations are in the treated group. So, $\alpha_{1}=1/2$, $\alpha_{2}=1/2$, and $\alpha_{0}=-1/4$. Following this, the sum of squared residuals $A$ from Equation (14) can be expressed as
$A=\sum_{i\in t=0,T}\left(1/4-1/2\right)^{2}+\sum_{i\in t=1,T}\left(1+1/4-1/2-1/2\right)^{2}+\sum_{i\in t=0,C}\left(-1/4\right)^{2}+\sum_{i\in t=1,C}\left(1/4-1/2\right)^{2}$
[…]
Using the same values, we can show that $\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)=\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)=\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)=\left(-\hat{\alpha}_{0}\right)=1/4$ from Equation (37). Therefore, we can show that $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)}{A}=\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}=\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}=\frac{\left(-\hat{\alpha}_{0}\right)}{A}=1/(N/4)$. With equal sample sizes, $N_{1,1}=N_{0,1}=N_{1,0}=N_{0,0}=N/4$.
Following this, we can conclude that $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group and the post-intervention period ($N_{1,1})$ and $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group in the pre-intervention period ($N_{1,0}$). Similarly, $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the control group in the post-intervention period $(N_{0,1})$, and $\frac{\left(-\hat{\alpha}_{0}\right)}{A}$ is reciprocal of the number of observations in the control group in the pre-intervention period ($N_{0,0})$. The same results hold for cases with unequal sample sizes.
With time-varying covariates, $\alpha_{3}$ is interpreted as $Pr(X_{i,t})=x$. Therefore, $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}-\hat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}+\hat{\alpha}_{3}x\right)}{G^{\prime}}$ are equivalent to the reciprocals of the number of observations in the treated group in the post-intervention period and the pre-intervention period, respectively, with $X_{i,t}=x$. Similarly, $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}+\hat{\alpha}_{3}x\right)}{G^{\prime}}$ and $\frac{\left(-\hat{\alpha}_{0}-\hat{\alpha}_{3}x\right)}{G^{\prime}}$ represent the probabilities for the observations being in the control group in the post- and pre-intervention periods, respectively, conditional on $X_{i,t}$.